OpenStudy (anonymous):

is y=1/3m+1 the same as y=x/3m+1?? if not then how do you find the equation of the straight line passing through the point (3,2) with y intercept 1?

OpenStudy (anonymous):

y = mx + b 2 = m(3) + 1 2 = 3m + 1 1 = 3m 1/3 = m So: y = (1/3)x + 1